OpenStudy (symone19):

4/10 ÷ 1/3 = 6/5 2/15 5/6 11/5 help me

OpenStudy (symone19):

9/5 - 4/3 = 1/3 8/15 7/15 1 1/15 this one two

OpenStudy (mathmale):

One problem at a time, please. Otherwise others won't know which problem you're focusing on at the moment. Which one do you want to work on?

OpenStudy (jiteshmeghwal9):

Hint :- \(\Large{\frac{a}{b}\div\frac{c}{d}=\frac{a}{b} \times \frac{d}{c}}\)

OpenStudy (symone19):

the first one